OpenStudy (anonymous):

What constant rate would be required to change the volume from the observed value at t=1.5 to the one at t=3? we call this average rate of change on the interval similar to the way average velocity is defined in this section). t= 0 .5 1 1.5 2 2.5 3 v=5 5.2 4.8 1.7 2.9 4.1 5

OpenStudy (amistre64):

its the slope of the points (1.5,1.7) and (3,5)

OpenStudy (amistre64):

y' = c y = cx + C

OpenStudy (amistre64):

find the equation to the line between the 2 points is about the best i can give you on it

OpenStudy (anonymous):

5-1.7/3-1.5?

OpenStudy (anonymous):

yea i cant find it. i keep getting different outcomes for each equation

OpenStudy (anonymous):

myininaya? could you help me?

OpenStudy (amistre64):

(3.0,5.0) -(1.5,1.7) ----------- 1.5, 3.3; slope = 33/15

OpenStudy (amistre64):

or 11/5 if your lucky lol

OpenStudy (amistre64):

y = mx -mPx + Py y = (11/5)x - 11(3)/5 + 5 y = (11/5)x -33/5 + 25/5 y = (11/5)x - 8/5 is what i get if its right

OpenStudy (anonymous):

but how does that get you the constant rate needed?

OpenStudy (amistre64):

"we call this average rate of change" the average rate of change is simply the slope of the line betweem the extremes on an interval

OpenStudy (amistre64):

your interval is from 1.5 to 3

OpenStudy (amistre64):

i believe the constant rate of change needed; would be just the slope itself

OpenStudy (anonymous):

ah yess 2.2 thank you!

OpenStudy (amistre64):



OpenStudy (anonymous):

could you help me with another rate of change?